# Enhancing Robot Navigation Policies with Task-Specific Uncertainty Management

### v2

@@ 4. Experiments
Figure 4. Comparison of navigation trajectories on the task: Start at the dock, navigate around the central fountain, then around the left fountain, and finally around the right fountain. The white line represents the ASV’s trajectory, with red points marking locations where it reduced uncertainty to meet task-specific requirements. Pink diamonds indicate collision points, the green rectangle marks the starting dock, and the TSUMs for GUIDEd agents are shown above their respective figures.
4.1. Experimental Setup
4.1.1. Environment
@@
4.1.2. Autonomous Surface Vehicle Setup
Our experiments utilize the Sea-Robotics Surveyor Autonomous Surface Vehicle (ASV), which is equipped with a Global Positioning System (GPS), an Inertial Measurement Unit (IMU), and a YSI EXO2 multiparameter water quality sonde. The GPS provides precise positioning, while the IMU assists with motion tracking and orientation. The YSI EXO2 sonde, primarily used for environmental monitoring, contributes to the overall state estimation by providing additional context.
@@ 4.1.2. Autonomous Surface Vehicle Setup
4.1.3. Position Estimation Modes
To represent the operational challenges of managing uncertainty during navigation, we model two modes of position estimation for the ASV (Puthumanaillam et al., 2024):
@@ 4.1.3. Position Estimation Modes
    Noisy Position Estimation: By default, the ASV estimates its position using the IMU and YSI EXO2 sensors, which results in a noisy and less accurate position estimate (Rojas et al., 2024). This mode represents the low-cost but high-uncertainty estimate.
@@
4.1.4. Evaluation Protocol
The agent receives a natural language task specification, describing a navigation task using key features of the environment [^2] Refer to Table 1 for details about task types. Task completion is evaluated based on specific criteria for each task type. For goal-reaching tasks, the agent must reach within a 1.5-meter radius of the target. For perimeter or exploration tasks, it must remain within a 3.5-meter corridor of the path, and for avoid tasks, a 3-meter minimum distance from designated areas is required. Task completion rate (TCR) is determined by the proportion of the task completed, and rewards are based on the total reward earned at task completion. All baselines interpret instructions through the same semantic parser, converting them into structured representations suitable for each algorithm.
Tasks Metric SAC SAC-P B-SAC CVaR RAA HEU G-PPO G-SAC Goal reaching: waypoint TCR (%) ($\uparrow$) 67.2% 82.1% 75.9% 68.8% 35.3% 71.3% 83.8% 95.1% visit [coordinate] Reward (avg) ($\uparrow$) 186.2 260.4 249.5 184.8 26.9 176.1 319.0 425.5 Goal reaching: context TCR (%) ($\uparrow$) 68.9% 84.3% 74.2% 66.8% 32.7% 73.5% 81.7% 92.0% navigate to dock Reward (avg) ($\uparrow$) 144.1 241.8 189.2 134.6 53.5 137.5 308.3 410.4 Avoid tasks TCR (%) ($\uparrow$) 71.3% 83.2% 79.6% 78.4% 51.3% 62.4% 89.6% 88.7% avoid the central fountain Reward (avg) ($\uparrow$) 177.8 199.2 277.6 220.4 107.8 174.4 437.6 482.2 Perimeter tasks TCR (%) ($\uparrow$) 44.3% 51.6% 56.3% 41.6% 39.8% 49.6% 74.9% 83.7% go around the left fountain Reward (avg) ($\uparrow$) 84.4 132.8 170.4 32.8 111.2 146.8 449.2 594.6 Explore tasks TCR (%) ($\uparrow$) 88.6% 92.4% 87.6% 84.9% 71.3% 88.7% 93.8% 97.5% explore top-right quadrant Reward (avg) ($\uparrow$) 486 474 526 449 402 537 548 595 Restricted areas TCR (%) ($\uparrow$) 70.8% 82.1% 78.4% 71.2% 51.1% 63.9% 90.0% 88.7% visit dock while avoiding top-right quadrant Reward (avg) ($\uparrow$) 266.4 306.8 307.2 269.6 201.4 261.2 570.0 634.6 Multi-goal tasks TCR (%) ($\uparrow$) 31.3% 42.9% 37.7% 30.9% 19.5% 42.1% 72.8% 81.7% Combination of tasks (see Fig. 3) Reward (avg) ($\uparrow$) 124.4 135.2 122.4 129.2 100.4 155.2 423.6 511.4
Table 1. Performance comparison between GUIDE and baseline methods on Task Completion Rate (TCR) and average cumulative reward, across different task categories. Results are averaged over 50 experiments per task category. Example tasks are provided below each task category to illustrate the types of navigation challenges evaluated.
4.2. Baselines and Ablations
@@ 4.2.1. Ablation Studies
    GUIDEd PPO (G-PPO, Ablation 2): We implement GUIDEd PPO (Schulman et al., 2017), incorporating TSUMs and agent uncertainty into the state representation, to examine the effect of the underlying RL algorithm on GUIDE’s performance.
4.2.2. Baselines
@@
    Bootstrapped Uncertainty-Aware RL (B-SAC): We implement Bootstrapped SAC (Kumar et al., 2019b), training multiple value networks to estimate epistemic uncertainty, which guides exploration and decision-making.
@@
    Risk-Aware RL (CVaR): We train agents using a CVaR-based RL algorithm (Kahn et al., 2017), optimizing a risk-sensitive objective to minimize potential high-cost outcomes, focusing on worst-case scenarios without task-specific uncertainty levels.
@@
    Uncertainty-Aware Motion Planning (RAA): We employ Risk-Aware A* (Hu and Jin, 2023) to plan paths that minimize collision probabilities under uncertainty, aiming for overall uncertainty minimization without considering task-specific requirements.
4.3. Results
The performance of GUIDE compared to various baseline methods is summarized in Table 1. Our results demonstrate that GUIDE consistently outperforms all baselines across all task categories. Notably, both GUIDEd Actor-Critic (G-SAC) and GUIDEd PPO (G-PPO) outperform the baselines; however, G-SAC generally achieves higher task completion rates and rewards.
4.3.1. Analysis of Ablation Studies
@@
To assess the significance of incorporating Task-Specific Uncertainty Maps (TSUMs), we compare the standard Soft Actor-Critic (SAC) without TSUMs to GUIDEd Actor-Critic (G-SAC). As shown in Table 1, conditioning on TSUMs significantly enhances performance across all tasks. Without TSUMs, SAC cannot manage positional uncertainty in a task-specific manner, leading to suboptimal navigation decisions and lower rewards. To further illustrate the effect of TSUM integration, we present in Figure 4 the trajectories of agents with and without TSUMs on a representative task. We observe that SAC often overuses high-precision localization in areas where it is unnecessary, incurring additional costs without significant benefits. Conversely, in critical regions requiring precise navigation, SAC fails to reduce uncertainty appropriately, leading to collisions with the fountains. In contrast, the G-SAC agent effectively uses TSUMs to adapt its uncertainty management, switching to high-precision localization in areas where the TSUM indicates low acceptable uncertainty. This enables the agent to navigate safely around obstacles and complete tasks more efficiently.
@@
We also investigate the impact of the underlying reinforcement learning algorithm by comparing GUIDEd PPO (G-PPO) to G-SAC, both of which integrate TSUMs but differ in their RL methodologies. This difference in the performance of G-SAC and G-PPO can be attributed to several factors inherent to the algorithms. G-SAC, based on the SAC framework, is an off-policy method that leverages entropy regularization to encourage exploration while maintaining stability. Its off-policy nature allows for more efficient sample utilization, which is particularly beneficial in continuous action spaces and when data collection is costly or limited. In contrast, PPO is an on-policy algorithm that relies on proximal updates to prevent large policy shifts, using a clipped objective function. While PPO is stable, it can be less sample-efficient (Liu, 2024), as it requires new data for each policy update and may not explore as effectively in complex environments.
Our empirical results and the trajectories in Figure 4 suggest that the off-policy efficiency and exploration capabilities of G-SAC make it better suited for navigation tasks. Furthermore, the entropy term in G-SAC encourages the agent to consider a wider range of actions, enabling it to discover more optimal strategies for managing uncertainty in a task-specific context.
@@
Figure 5. Comparison of navigation trajectories between GUIDEd SAC and SAC-P agents on the task: Start and end at the dock. Go around the perimeter of the area and visit coordinates [redacted for paper anonymity]. The light blue line represents the trajectory taken by the agent. Red dots indicate locations where the agent actively reduced its state estimation uncertainty. The green rectangle denotes the dock, blue rectangle marks the specific coordinates the agent is instructed to visit.
4.3.2. Comparison with Baselines
@@
GUIDE consistently outperforms all baselines across task types in both task completion rate and average reward (see Table 1). Standard RL methods like SAC and SAC-P lack task-specific uncertainty management. SAC-P penalizes high uncertainty uniformly, resulting in overly conservative behavior where precision is unnecessary and insufficient caution in critical regions, ultimately leading to lower performance. B-SAC estimates epistemic uncertainty but fails to adapt to task-specific requirements, leading to inefficient exploration. Risk-Aware RL methods like CVaR are uniformly risk-averse, missing opportunities for calculated risk-taking that could improve task success. RAA aims to minimize overall uncertainty without considering task context, often generating inefficient paths. The Heuristic Policy (HEU) switches to exact localization near obstacles, but lacks GUIDE’s adaptability, failing to adjust to sudden changes in uncertainty requirements.
4.3.3. Behavior of GUIDEd Agents
@@
We observed distinct behaviors in GUIDEd agents across various tasks. Analyzing the specific task shown in Figure 5, GUIDEd agents strategically adjust their reliance on precise position estimation versus noisier estimates. In areas where the TSUMs indicate high precision is necessary — such as navigating near obstacles or close to the perimeters (highlighted in yellow) — the agents opt for exact positioning despite the higher operational cost. Conversely, in less critical regions (shown in purple), they rely on less precise, noisy estimates. This adaptability allows GUIDEd agents to manage uncertainty more efficiently than baselines, resulting in faster completion times and smoother trajectories. Although not perfect — occasionally missing sections of the perimeter (indicated by black shaded regions in Figure 5) — GUIDEd agents significantly outperform baselines like SAC-P with engineered rewards. Baseline methods, lacking adaptive uncertainty management, frequently fail to complete tasks safely. They often collide with obstacles (Figure 5 pink diamond ) or take inefficient paths (Figure 5 black region).
5. Conclusion
@@
In this paper, we introduced GUIDE, a framework that integrates task-specific uncertainty requirements into robotic navigation policies. Central to our approach is the concept of Task-Specific Uncertainty Maps (TSUMs), which represent acceptable levels of state estimation uncertainty across different regions of the environment based on the given task. By conditioning navigation policies on TSUMs, we enable robots to reason about the context-dependent importance of certainty and adapt their behavior accordingly.
@@
We demonstrated how GUIDE can be incorporated into reinforcement learning frameworks by augmenting the state representation to include both the acceptable uncertainty levels from TSUMs and the robot’s current state estimation uncertainty. Specifically, we adapted the Soft Actor-Critic (SAC) algorithm to operate in this augmented state space, resulting in the GUIDEd SAC algorithm. Our extensive experiments on a variety of real-world robotic navigation tasks showed that policies conditioned on TSUMs significantly improve task completion rates and overall performance compared to baseline methods that do not explicitly consider task-specific uncertainty. The results indicate that incorporating task-specific uncertainty requirements allows robots to effectively balance task objectives with appropriate uncertainty management. GUIDE enhances navigation strategies by enabling robots to focus their uncertainty reduction efforts where it is most critical for task success, leading to improved efficiency.
GUIDE is designed to be flexible and can be applied to various robotic platforms and tasks. Future work includes extending the framework to handle dynamic environments where uncertainty requirements may change over time.
SUPPLEMENTARY MATERIAL
@@
Appendix A Additional Experimental Results
(a) "Go to point [80,90] and then go around the central fountain and return to the dock."
(b) "Go to [80,60] and then go to [80, 20] and then go to [80, 70] and finally return to the dock. This task has to be completed by avoiding the right half of the area."
(c) "Go to [40,60] while avoiding the left half of the environment."
(d) "Explore the top half of the environment."
(e) "Navigate the perimeter of the bottom half of the lake."
(f) "Go to [40,60]."
Figure 6. Illustrations of various navigation tasks performed by GUIDEd SAC in the environment. In all images, the light blue line represents the trajectory, red dots indicate areas where uncertainty was reduced, the green rectangle represents the dock, and brown rectangles represent waypoints that need to be reached. The specific tasks are described below each figure.

### v3

@@ IV Experiments
Figure 4: Comparison of navigation trajectories on the task: Start at the dock, navigate around the central fountain, then around the left fountain, and finally around the right fountain. The white line represents the ASV’s trajectory, with red points marking locations where it reduced uncertainty to meet task-specific requirements. Pink diamonds indicate collision points, the green rectangle marks the starting dock, and the TSUMs for GUIDEd agents are shown above their respective figures.
IV-A Experimental Setup
IV-A1 Environment
@@
IV-A2 Autonomous Surface Vehicle Setup
Our experiments utilize the SeaRobotics Surveyor Autonomous Surface Vehicle (ASV), which is equipped with a Global Positioning System (GPS), an Inertial Measurement Unit (IMU), and a YSI EXO2 multiparameter water quality sonde. The GPS provides precise positioning, while the IMU assists with motion tracking and orientation. The YSI EXO2 sonde, primarily used for environmental monitoring, contributes to the overall state estimation by providing additional context.
@@ IV-A2 Autonomous Surface Vehicle Setup
IV-A3 Position Estimation Modes
To represent the operational challenges of managing uncertainty during navigation, we model two modes of position estimation for the ASV [49]:
@@ IV-A3 Position Estimation Modes
    Noisy Position Estimation: By default, the ASV estimates its position using the IMU and YSI EXO2 sensors, which results in a noisy and less accurate position estimate [53]. This mode represents the low-cost but high-uncertainty estimate.
@@
IV-A4 Evaluation Protocol
The agent receives a natural language task specification, describing a navigation task using key features of the environment. Task completion is evaluated based on specific criteria for each task type. For goal-reaching tasks, the agent must reach within a 1.5-meter radius of the target. For perimeter or exploration tasks, it must remain within a 3.5-meter corridor of the path, and for avoid tasks, a 3-meter minimum distance from designated areas is required. Task completion rate (TCR) is determined by the proportion of the task completed, and rewards are based on the total reward earned at task completion. All baselines interpret instructions through the same semantic parser, converting them into structured representations suitable for each algorithm.
Tasks Metric SAC SAC-P B-SAC CVaR RAA HEU G-PPO G-SAC Goal reaching: waypoint TCR (%) ($\uparrow$) 67.2% 82.1% 75.9% 68.8% 35.3% 71.3% 83.8% 95.7% visit [coordinate] Reward (avg) ($\uparrow$) 186.2 260.4 249.5 184.8 26.9 176.1 319.0 429.2 Goal reaching: context TCR (%) ($\uparrow$) 68.9% 84.3% 74.2% 66.8% 32.7% 73.5% 81.7% 91.5% navigate to dock Reward (avg) ($\uparrow$) 144.1 241.8 189.2 134.6 53.5 137.5 308.3 400.2 Avoid tasks TCR (%) ($\uparrow$) 71.3% 83.2% 79.6% 78.4% 51.3% 62.4% 89.6% 87.4% avoid the central fountain Reward (avg) ($\uparrow$) 177.8 199.2 277.6 220.4 107.8 174.4 437.6 480.2 Perimeter tasks TCR (%) ($\uparrow$) 44.3% 51.6% 56.3% 41.6% 39.8% 49.6% 74.9% 85.6% go around the left fountain Reward (avg) ($\uparrow$) 84.4 132.8 170.4 32.8 111.2 146.8 449.2 599.6 Explore tasks TCR (%) ($\uparrow$) 88.6% 92.4% 87.6% 84.9% 71.3% 88.7% 93.8% 97.3% explore top-right quadrant Reward (avg) ($\uparrow$) 486 474 526 449 402 537 548 595 Restricted areas TCR (%) ($\uparrow$) 70.8% 82.1% 78.4% 71.2% 51.1% 63.9% 90.0% 88.6% visit dock while avoiding top-right quadrant Reward (avg) ($\uparrow$) 266.4 306.8 307.2 269.6 201.4 261.2 570.0 634.6 Multi-goal tasks TCR (%) ($\uparrow$) 31.3% 42.9% 37.7% 30.9% 19.5% 42.1% 72.8% 81.7% Combination of tasks (see Fig. 3) Reward (avg) ($\uparrow$) 124.4 135.2 122.4 129.2 100.4 155.2 423.6 510.4
TABLE I: Performance comparison between GUIDE and baseline methods on Task Completion Rate (TCR) and average cumulative reward, across different task categories. Results are averaged over 50 experiments per task category. Example tasks are provided below each task category to illustrate the types of navigation challenges evaluated.
IV-B Baselines and Ablations
@@ IV-B1 Ablation Studies
    GUIDEd PPO (G-PPO, Ablation 2): We implement GUIDEd PPO [56], incorporating TSUMs and agent uncertainty into the state representation, to examine the effect of the underlying RL algorithm on GUIDE’s performance.
IV-B2 Baselines
@@
    Bootstrapped Uncertainty-Aware RL (B-SAC): We implement Bootstrapped SAC [29], training multiple value networks to estimate epistemic uncertainty, which guides exploration and decision-making.
@@
    Risk-Aware RL (CVaR): We train agents using a CVaR-based RL algorithm [24], optimizing a risk-sensitive objective to minimize potential high-cost outcomes, focusing on worst-case scenarios without task-specific uncertainty levels.
@@
    Uncertainty-Aware Motion Planning (RAA): We employ Risk-Aware A* [20] to plan paths that minimize collision probabilities under uncertainty, aiming for overall uncertainty minimization without considering task-specific requirements.
IV-C Results
The performance of GUIDE compared to various baseline methods is summarized in Table I [^3] Extensive additional experiment results can be found in the appendix and the video supplementary.. Our results demonstrate that GUIDE consistently outperforms all baselines across all task categories. Notably, both GUIDEd Actor-Critic (G-SAC) and GUIDEd PPO (G-PPO) outperform the baselines; however, G-SAC generally achieves higher task completion rates and rewards.
IV-C1 Analysis of Ablation Studies
@@
To assess the significance of incorporating Task-Specific Uncertainty Maps (TSUMs), we compare the standard Soft Actor-Critic (SAC) without TSUMs to GUIDEd Actor-Critic (G-SAC). As shown in Table I, conditioning on TSUMs significantly enhances performance across all tasks. Without TSUMs, SAC cannot manage positional uncertainty in a task-specific manner, leading to suboptimal navigation decisions and lower rewards. To further illustrate the effect of TSUM integration, we present in Fig. 4 the trajectories of agents with and without TSUMs on a representative task. We observe that SAC often overuses high-precision localization in areas where it is unnecessary, incurring additional costs without significant benefits. Conversely, in critical regions requiring precise navigation, SAC fails to reduce uncertainty appropriately, leading to collisions with the fountains. In contrast, the G-SAC agent effectively uses TSUMs to adapt its uncertainty management, switching to high-precision localization in areas where the TSUM indicates low acceptable uncertainty. This enables the agent to navigate safely around obstacles and complete tasks more efficiently.
@@
We also investigate the impact of the underlying reinforcement learning algorithm by comparing GUIDEd PPO (G-PPO) to G-SAC, both of which integrate TSUMs but differ in their RL methodologies. This difference in the performance of G-SAC and G-PPO can be attributed to several factors inherent to the algorithms. G-SAC, based on the SAC framework, is an off-policy method that leverages entropy regularization to encourage exploration while maintaining stability. Its off-policy nature allows for more efficient sample utilization, which is particularly beneficial in continuous action spaces and when data collection is costly or limited. In contrast, PPO is an on-policy algorithm that relies on proximal updates to prevent large policy shifts, using a clipped objective function. While PPO is stable, it can be less sample-efficient [34], as it requires new data for each policy update and may not explore as effectively in complex environments.
Our empirical results suggest that the off-policy efficiency and exploration capabilities of G-SAC make it better suited for navigation tasks. Furthermore, the entropy term in G-SAC encourages the agent to consider a wider range of actions, enabling it to discover more optimal strategies for managing uncertainty in a task-specific context.
@@
Figure 5: Comparison of navigation trajectories between GUIDEd SAC and SAC-P agents on the task: Start and end at the dock. Go around the perimeter of the area and visit coordinates [redacted for paper anonymity]. The light blue line represents the trajectory taken by the agent. Red dots indicate locations where the agent actively reduced its state estimation uncertainty. The green rectangle denotes the dock, blue rectangle marks the specific coordinates the agent is instructed to visit.
IV-C2 Comparison with Baselines
@@
GUIDE consistently outperforms all baselines across task types in both task completion rate and average reward (see Table I). Standard RL methods like SAC and SAC-P lack task-specific uncertainty management. SAC-P penalizes high uncertainty uniformly, resulting in overly conservative behavior where precision is unnecessary and insufficient caution in critical regions, ultimately leading to lower performance. B-SAC estimates epistemic uncertainty but fails to adapt to task-specific requirements, leading to inefficient exploration. Risk-Aware RL methods like CVaR are uniformly risk-averse, missing opportunities for calculated risk-taking that could improve task success. RAA aims to minimize overall uncertainty without considering task context, often generating inefficient paths. The Heuristic Policy (HEU) switches to exact localization near obstacles, but lacks GUIDE’s adaptability, failing to adjust to sudden changes in uncertainty requirements.
IV-C3 Behavior of GUIDEd Agents
@@
We observed distinct behaviors in GUIDEd agents across various tasks. Analyzing the specific task shown in Fig. 5, GUIDEd agents strategically adjust their reliance on precise position estimation versus noisier estimates. In areas where the TSUMs indicate high precision is necessary — such as navigating near obstacles or close to the perimeters (Fig. 5, highlighted in yellow) — the agents opt for exact positioning despite the higher operational cost. Conversely, in less critical regions (Fig. 5, highlighted in purple), they rely on less precise, noisy estimates. This adaptability allows GUIDEd agents to manage uncertainty more efficiently than baselines, resulting in faster completion times and smoother trajectories. Although not perfect — occasionally missing sections of the perimeter (indicated by black shaded regions in Fig. 5) — GUIDEd agents significantly outperform baselines like SAC-P with engineered rewards. Baseline methods, lacking adaptive uncertainty management, frequently fail to complete tasks safely. They often collide with obstacles (pink diamond highlighted in Fig. 5) or take inefficient paths (Fig. 5 black highlighted region).
V Limitations
While GUIDE demonstrates promising results, several limitations should be noted. First, the current implementation assumes static uncertainty requirements and may not adapt well to environments where acceptable uncertainty levels change rapidly. Second, the framework relies heavily on the quality of image and may face challenges in environments with bad-quality visual data. Third, while our approach effectively handles a wide range of navigation tasks, it may require additional consideration for scenarios where task specifications are inherently ambiguous or contain multiple valid interpretations of uncertainty requirements.
VI Conclusion
@@
In this paper, we introduced GUIDE, a framework that integrates task-specific uncertainty requirements into robotic navigation policies for limited-localization environments. Central to our approach is the concept of Task-Specific Uncertainty Maps (TSUMs), which represent acceptable levels of state estimation uncertainty across different regions of the environment based on the given task. By conditioning navigation policies on TSUMs, we enable robots to reason about the context-dependent importance of certainty and adapt their behavior accordingly.
@@
We demonstrated how GUIDE can be incorporated into reinforcement learning frameworks by augmenting the state representation to include both the acceptable uncertainty levels from TSUMs and the robot’s current state estimation uncertainty. Specifically, we adapted the Soft Actor-Critic (SAC) algorithm to operate in this augmented state space, resulting in the GUIDEd SAC algorithm. Our extensive experiments demonstrate GUIDE’s effectiveness through successful in-the-wild deployments across various previously unseen environments, showing significant improvements in task completion rates and overall performance compared to baseline methods that do not explicitly consider task-specific uncertainty.
The results indicate that incorporating task-specific uncertainty requirements allows robots to effectively balance task objectives with appropriate uncertainty management. GUIDE enhances navigation strategies by enabling robots to focus their uncertainty reduction efforts where it is most critical for task success, leading to improved efficiency. Future work includes extending the framework to handle dynamic environments where uncertainty requirements may change over time.
Appendix
@@
  • PFA Video Demonstrations: Selected experimental runs are included in the accompanying video.
VII Additional Experimental Results
((a)) "Go to point [80,90] and then go around the central fountain and return to the dock."
((b)) "Go to [80,60] and then go to [80, 20] and then go to [80, 70] and finally return to the dock. This task has to be completed by avoiding the right half of the area."
((c)) "Go to [40,60] while avoiding the left half of the environment."
((d)) "Explore the top half of the environment."
((e)) "Navigate the perimeter of the bottom half of the lake."
((f)) "Go to [40,60]."
Figure 6: Illustrations of various navigation tasks performed by GUIDEd SAC in the environment. In all images, the light blue line represents the trajectory, red dots indicate areas where uncertainty was reduced, the green rectangle represents the dock, and brown rectangles represent waypoints that need to be reached. The specific tasks are described below each figure.